OpenStudy (anonymous):

Any help would be great: Function Notation If ƒ(x ) = x 2 + 1 and g(x ) = 3x + 1, find 2 · ƒ(4).

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

Hi

OpenStudy (misty1212):

are you looking for \[2\times f(4)\]?

OpenStudy (anonymous):

Those values are subbed into the expression

OpenStudy (anonymous):

*equation

OpenStudy (misty1212):

there is not \(g\) in this your main job is to find \[f(4)=4^2+1\]

OpenStudy (misty1212):

then multiply the result by 2

OpenStudy (misty1212):

you got that?

OpenStudy (misty1212):

\[f(x)=^2+1\\ f(4)=4^2+1=17\\ 2f(4)=2\times 17\]

OpenStudy (anonymous):

So 34. Why would you not add g of x?

OpenStudy (misty1212):

you asked for \[2f(4)\]right?

OpenStudy (misty1212):

got no \(g\) in it

OpenStudy (anonymous):

That is where I reached my snag there was no x value so I assumed you simplified.

OpenStudy (misty1212):

lol there is no such thing as "simplify"

OpenStudy (misty1212):

add, subtract, multiply, divide, combine like terms, reduce to lowest terms, write in simplest radical form etc. but not "simplify"

OpenStudy (anonymous):

Great so I just have to solve for F(X)?

OpenStudy (anonymous):

My original question was just what was on my paper.

OpenStudy (anonymous):

So another example. If ƒ(x ) = x 2 + 1 and g(x ) = 3x + 1, find [ƒ(4)] 2. Would just be 257?

OpenStudy (anonymous):

x^2 and [f(4)] ^ 2

OpenStudy (beginnersmind):

well f(4) is 17. So [f(4)]^2 is 17^2.

OpenStudy (beginnersmind):

Or did you change the definition of f(x) to f(x) = x^2? Then [f(4)]^2 = (4^2)^2 = 16^2 = 256